(COGS and COGM: Two products) In its second year of operations, Memories, Inc. has decided to expand the product e by producing replicas of historic buildings. These replicas will require the purchase of new building molds at a cost of $.18 per replica. Of course, new doll molds will not be equired. All other materials and prices will remain the same. The replicas require additional processing time because of the details on the uildings that limits production to 18 replicas per hour per assembly line. The replicas re not expected to affect the sales of dolls. In the second year of operations, MI xpects to produce and sell about 352,800 dolls and 25,200 replicas. increasing production is expected to increase overhead costs by 5% in Year 2. Direct labor costs per hour are not expected to change, but the number of labor hours is stimated to be 94,500. The costs of product promotion and advertising are expected to crease to $3,000 per month. All other selling and administrative costs are expected to emain the same as in Year 1 Actual production in Year 2 was 345,132 dolls and 25,200 replicas. Direct material osts transferred to Work-in-Process were $259,000 for dolls and $15,624 for replicas. irect labor costs were $862,830 for the dolls and $70,010 for replicas, representing 86,283 and 7,001 direct labor hours, respectively. Actual overhead costs were 203,600
media%2F2ab%2F2ab7404f-6f71-43f7-88ba-6f
Answer #1

Answer - A

Planned units of production for two products should be the basis of overhead allocation since the processing time for both is different and there is an impact on the overall overhead cost of 5% increase because of new product.

Answer - B

O/H share for DOLLS = 203600/(352800+25200)*352800 = 190,027

Rate per unit for DOLLS = 190027 / 352800 = 0.54

O/H share for REPLICA = 203600/(352800+25200)*25200 = 13,573

Rate per unit for REPLICA = 13573 / 25200 = 0.54

Answer - C

Cost of Production is as hereunder

Costing Sheet Details Units Produced DOLLS REPLICAS 345132 25200 25900015 Direct Material Cost Direct Labour Cost O/H share @

Total O/H absorbed = 186371 + 13608 = 199979

Hence underabsorption of = 203600 - 199979 = 3621